what equations solve the problem: what 2 digit number multiplied by 2 and adding 2 has it's digits reversed?
Answer by richard1234(7193)

Suppose the number is 10a + b, where a,b are the digits

2(10a+b) = 10b+a + 2

20a + 2b = 10b + a + 2

19a - 8b = 2

Then solve for integer solutions a,b (between 0 - 9).